### v1

Abstract
Real-World Assets (RWAs) have recently attracted increasing attention as a means of bridging traditional financial instruments with decentralized infrastructures. By representing assets such as bonds, commodities, and real estate on blockchains, RWAs can enhance liquidity, broaden accessibility, and extend the scope of decentralized finance. Industry forecasts further suggest rapid growth of tokenized RWAs in the coming years, underscoring their potential role in the evolution of digital financial markets. However, when deployed across multiple blockchains, RWAs face challenges such as repeated authentication on different chains and inefficiency caused by multi-step settlement protocols. To address these issues, we present a cross-chain framework for RWAs that emphasizes identity management, authentication, and interaction. The framework integrates Decentralized Identifiers and Verifiable Credentials with customized attributes to support decentralized identification, and incorporates an authentication protocol based on Simplified Payment Verification to avoid redundant verification across chains. Furthermore, we design a cross-chain channel that enables the settlement of RWAs without requiring channel closure, thereby improving operational efficiency. We implement the framework and evaluate it through simulations, which confirm its feasibility and demonstrate improvements in efficiency for RWAs in cross-chain settings.
keywords:
[…]
1 Introduction
Real-World Assets (RWAs) refer to the tokenization of off-chain assets such as government bonds, commodities, and real estate, enabling them to be represented and transacted on blockchain systems [5]. They are attracting increasing attention in Web3 finance: as of June 2025, the total market capitalization of global stablecoins is approximately \$250 billion, while the market size of RWA tokenization grew from \$8.6 billion at the beginning of 2025 to \$23 billion in the first half of the year, representing a 260% increase. Looking ahead, Citibank estimates that tokenized RWAs may reach a scale of around \$4 trillion by 2030, indicating their potential role in the future development of digital financial markets [7, 23].
Blockchain [28, 11, 9] constitutes the key infrastructure for the tokenization and circulation of RWAs, providing transparency and immutability for asset issuance and transfer. In recent years, blockchain systems have expanded rapidly. By June 2025, CoinMarketCap reported more than ten thousand active cryptocurrency platforms [6], suggesting that RWAs can be deployed across an increasing number of heterogeneous blockchains. Although such distribution facilitates broader adoption [17, 11, 22], the confinement of RWAs to single blockchains limits their financial utility, as assets cannot be seamlessly transferred or traded across different chains. This situation is analogous to traditional financial markets where banks are unable to conduct transactions across borders, thereby restricting the mobility and efficiency of capital. Fortunately, cross-chain technologies [2, 13, 12], widely regarded as a primary means of overcoming information silos between blockchains, may provide a feasible approach to enabling RWA interactions across heterogeneous systems. Nevertheless, existing research and practice on RWAs remain confined to single-chain contexts and have not systematically examined their cross-chain realization (shown in Figure 1). To address this limitation, we propose a cross-chain RWA framework that aims to resolve the main challenges inherent in cross-chain processes.
Figure 1: The multi-chain environment creates opportunities for the cross-chain circulation of RWAs. However, redundant authentication and inefficient cross-chain operations remain critical challenges.
Cross-Chain Redundant Authentication. Compared with single-chain deployment, cross-chain scenarios require an RWA to undergo repeated authentications during inter-chain transfers. In a single-chain setting, an RWA only needs to be authenticated once within the local ledger, and its validity can then be consistently recognized across applications on that chain. In cross-chain environments, each blockchain maintains independent trust assumptions and verification rules. As a result, existing approaches typically require RWAs to undergo full authentication each time they are transferred on a different chain. Ideally, an RWA authenticated on one chain should be recognized by another through a proof of its prior authentication, thereby avoiding redundant procedures. Such a mechanism is particularly relevant when assets move from highly trusted chains, such as Ethereum mainnet, to less secure environments like its sidechains.
Inefficiency in Cross-Chain Operations. Even within a single blockchain, operations involving RWAs already incur latency due to consensus. For instance, the two most prominent blockchain platforms, Bitcoin and Ethereum, support approximately 7 TPS and 15 TPS, which is significantly lower than traditional payment networks such as Visa, capable of processing several thousand TPS [19]. In cross-chain environments, this inefficiency is further amplified, since completing a cross-chain interaction typically requires multiple interdependent operations across heterogeneous ledgers. To ensure atomicity, decentralized protocols such as Hashed Timelock Contracts (HTLCs) are commonly employed; however, they generally involve four distinct on-chain operations, which substantially prolong completion time and increase operational costs. Cross-chain channels [13, 15] attempt to mitigate this limitation by integrating HTLC-based mechanisms with off-chain channels to improve efficiency. However, in existing designs, each settlement requires closing the channel and reopening a new one for subsequent interactions, which introduces additional costs and delays. Therefore, the central challenge is to design mechanisms that avoid repeated channel closure during settlement, thereby enabling more efficient cross-chain interoperability for RWAs.
To fill this gap, we propose a cross-chain framework for RWAs to redefine the workflow of RWAs, with particular emphasis on identification, cross-chain authentication, and cross-chain interaction. To ensure decentralization throughout the process, we adopt Decentralized Identifiers (DIDs) and Verifiable Credentials (VCs), which are widely used in Web3, and extend them with customized attributes tailored to represent RWAs. Building on Simplified Payment Verification (SPV), we design a cross-chain authentication protocol that prevents redundant verification of RWAs across multiple chains. Finally, we introduce a novel cross-chain channel scheme that enables the settlement of RWAs without requiring channel closure, thereby supporting more efficient cross-chain interactions.
For convenience, we highlight our main contributions as follows:
  1. 1.
    We propose a cross-chain framework for RWAs that considers authentication and efficiency in cross-chain processes. To the best of our knowledge, this is the first work in this direction.
  2. 2.
    To address the problem of redundant authentication across multiple chains, we design a cross-chain authentication protocol that enables RWAs authenticated on one chain to be recognized by others.
  3. 3.
    To improve efficiency in cross-chain operations, we develop a novel channel scheme that supports the settlement of RWAs without requiring channel closure.
  4. 4.
    We implement and conduct simulation experiments to demonstrate its feasibility and evaluate its performance.
The rest of the paper is organized as follows. In Section 2, we review the most related work. Section 3 introduces the necessary preliminary knowledge. In Section 4, we present our framework and discuss its applications. Section 5 reports simulation experiments that evaluate the performance of our scheme. Finally, Section 6 concludes the paper with remarks and insights.
2 Related Works
In this section, we review related work on Real-World Assets (RWAs) from both industry and academia.
2.1 RWAs in Industry
RWAs have already been actively deployed across various industry sectors. MakerDAO [8] integrated tokenized U.S. Treasuries into its collateral framework, reinforcing the stability of decentralized finance. HSBC [1] introduced a gold-backed RWA token for retail investors in Hong Kong. Securitize [21] operates a regulated platform for issuing and trading RWAs, enabling compliant digital access to private companies and funds. Brickken [4] has tokenized over USD 250 million in RWAs across 14 countries. Goldman Sachs and BNY Mellon [3] partnered to tokenize money market fund shares on blockchain, aiming to improve accessibility and settlement efficiency.
2.2 RWAs in Academia
In academia, research on real-world assets has primarily focused on single-chain architectures and survey studies. Chen et al. [5] analyzed the security of RWAs, highlighting risks in KYC/AML processes, oracle designs, and collateral structures. The work in [16] investigated the connection between RWAs and escrow mechanisms and discussed potential directions for future development. Zhao et al. [29] proposed a scalable framework that integrates Ethereum staking, Layer-2 rollups, and DAO governance for RWA management. Xia et al. [26] studied tokenization approaches for RWAs with the aim of enhancing liquidity and improving asset management practices. LEGO [18] provided a quantitative methodology that maps historical failures in blockchain systems to current designs, including those involving RWAs.
2.3 Summary and Motivations
Recent developments indicate that RWAs have already been deployed across multiple sectors of the financial industry, with initiatives ranging from tokenized government bonds and money market funds to gold and real estate. At the same time, academic research has largely concentrated on single-chain architectures and survey studies. This contrast reveals a critical gap: although RWAs are increasingly deployed on different blockchain systems, they remain isolated within their respective environments, lacking mechanisms for interoperability or cross-chain recognition. Each implementation therefore operates in a silo, limiting the broader financial utility of tokenized assets. Motivated by this observation, we propose a cross-chain framework for RWAs that addresses the challenges of redundant authentication and low efficiency, and enables secure, decentralized interoperability across multiple blockchains.
3 Preliminaries
In this section, we present decentralized identifiers, verifiable credentials, simple payment verification, off-chain channels, and hashed timelock contracts, which serve as the building blocks of our work.
3.1 Decentralized Identifiers
Decentralized Identifiers (DIDs) provide a decentralized identity mechanism for Web3, enabling users to prove ownership and control of their identities without relying on centralized authorities [24].
Definition 1 (DID Operations)
The process of a DID is represented by a tuple of polynomial-time algorithms $\text{DID}\overset{\text{def}}{=}(\mathsf{KeyGen},\mathsf{Create},\mathsf{Resolve},\mathsf{Update},\mathsf{Deactivate})$:
  • $(pk,sk)\leftarrow\mathsf{KeyGen}(\mathsf{pp})$. Given public parameters $\mathsf{pp}$, the algorithm generates a key pair $(pk,sk)$.
  • $(did,doc)\leftarrow\mathsf{Create}(sk,pk)$. Given a key pair $(sk,pk)$, the algorithm outputs a new identifier diddid and its associated document docdoc.
  • $doc\leftarrow\mathsf{Resolve}(did)$. Given an identifier diddid, the algorithm returns the current document docdoc linked to diddid.
  • $doc^{\prime}\leftarrow\mathsf{Update}(did,doc)$. Given an identifier diddid and its document docdoc, the algorithm produces an updated document $doc^{\prime}$, reflecting revised verification methods or service endpoints under the controller’s authorization.
  • $\varnothing\leftarrow\mathsf{Deactivate}(did)$. Given an identifier diddid, the algorithm deactivates it, making the identifier invalid and non-resolvable.
Decentralized Identifiers have been widely adopted in Web3 ecosystems and are formally standardized by the World Wide Web Consortium (W3C) [20].
3.2 Verifiable Credentials
Verifiable Credentials (VCs) provide issuer-signed statements about a DID subject that can be verified by cryptographic proofs without trusting the presenter [25].
Definition 2 (VC Operations)
Let the user be $\mathcal{U}$ with keys $(pk_{\mathcal{U}},sk_{\mathcal{U}})$ and DID $did_{\mathcal{U}}$, the issuer be $\mathcal{I}$ with $(pk_{\mathcal{I}},sk_{\mathcal{I}})$ and DID $did_{\mathcal{I}}$, and the verifier be $\mathcal{V}$. Let $\mathcal{R}^{\mathcal{I}}$ denote the issuer’s revocation list. The process is represented by the algorithm tuple $\Pi\overset{\text{def}}{=}(Req,\mathsf{Issue},\mathsf{Prove},\mathsf{Verify},\mathsf{Revoke})$:
  • $req\leftarrow\mathsf{Req}(\{d_{i}\}_{\mathcal{U}},sk_{\mathcal{U}})$. Given data items $\{d_{i}\}$, $\mathcal{U}$ signs a request reqreq to obtain a credential.
  • $cred\leftarrow\mathsf{Issue}(req,sk_{\mathcal{I}})$. Given reqreq, $\mathcal{I}$ signs and returns a credential credcred.
  • $\pi\leftarrow\mathsf{Prove}(cred,sk_{\mathcal{U}},\phi)$. Given credcred and a predicate $\phi$, $\mathcal{U}$ produces a proof $\pi$.
  • $\{0,1\}\leftarrow\mathsf{Verify}(\mathcal{R}^{\mathcal{I}},cred,\phi,\pi,did_{\mathcal{I}})$. $\mathcal{V}$ validates $\pi$, checks revocation against $\mathcal{R}^{\mathcal{I}}$, and confirms $did_{\mathcal{I}}$.
  • $\mathcal{R}^{\mathcal{I}\,^{\prime}}\leftarrow\mathsf{Revoke}(\mathcal{R}^{\mathcal{I}},cred,sk_{\mathcal{I}})$. $\mathcal{I}$ updates the status list to revoke or suspend credcred.
An important feature of VCs is selective disclosure, which allows a holder to reveal only specific attributes of a credential while keeping all other information hidden, thereby preserving privacy during verification [27].
3.3 Simple Payment Verification
SPV (Simple Payment Verification) enables a verifier to check whether a transaction is included in the blockchain without downloading full blocks [28].
Definition 3 (SPV)
The SPV process is represented by a tuple of polynomial-time algorithms $\text{SPV}\overset{\text{def}}{=}(\mathsf{Prove},\mathsf{Verify})$.
  • $\pi\leftarrow\mathsf{Prove}(tx,B)$. Given a transaction $tx$ and a block $B$ containing $tx$, the algorithm outputs a Merkle proof $\pi$, which consists of a Merkle path $p$ authenticating $tx$ with respect to the Merkle root $\mu$ in the header of $B$.
  • $\{0,1\}\leftarrow\mathsf{Verify}(tx,\pi,\mu)$. Given a transaction $tx$, a Merkle proof $\pi=(p,\mu)$, and a Merkle root $\mu$, the algorithm recomputes the authentication path using $p$ and returns 1 if and only if $tx$ is correctly included in the block, and 0 otherwise.
The Merkle path $p$ denotes the sequence of sibling hashes along the path from the transaction node to the root of the Merkle tree, and $\mu$ is the Merkle root contained in the corresponding block header.
3.4 Off-Chain Channel
An off-chain channel allows two parties to perform frequent transactions without recording each interaction on the blockchain, thereby reducing on-chain overhead while maintaining security guarantees through cryptographic commitments [10].
Definition 4 (Off-Chain Channel)
The off-chain channel process can be represented by a tuple of polynomial-time algorithms $\Pi\overset{\text{def}}{=}(\mathsf{Open},\mathsf{Update},\mathsf{Close})$.
  • $\mathsf{Open}$: Both parties deposit assets into a contract on-chain, which locks funds as collateral for subsequent off-chain interactions.
  • $\mathsf{Update}$: The parties exchange signed messages off-chain to update balances or states. Each new state invalidates the previous one and is enforceable on-chain in case of disputes.
  • $\mathsf{Close}$: When the channel is terminated, the latest signed state is submitted to the contract, and the locked deposits are returned to the parties’ accounts according to this final state, concluding the channel lifecycle.
Off-chain channels significantly improve scalability by shifting frequent interactions off-chain, while preserving security through cryptographic commitments and enforceability on-chain.
3.5 Hashed Timelock Contracts
Hashed Timelock Contracts (HTLCs) are conditional payment contracts widely used in cross-chain atomic swaps and payment channel networks [13]. They ensure that either both parties complete the transaction successfully or both are refunded, preventing unilateral loss.
Definition 5 (HTLC)
HTLC can be represented by a tuple of polynomial-time algorithms $\text{HTLC}\overset{\text{def}}{=}(\mathsf{Lock},\mathsf{Unlock},\allowbreak\mathsf{Refund})$.
  • $\mathsf{Lock}$: The sender generates a random preimage $\rho$ and computes its hash $h(\rho)$. The asset is locked in a contract with condition $h(\rho)$ and a timeout $T$.
  • $\mathsf{Unlock}$: The receiver obtains $\rho$ before the timeout and submits it to the contract, which verifies $h(\rho)$ and transfers the locked asset to the receiver.
  • $\mathsf{Refund}$: If the receiver fails to provide $\rho$ before the timeout, the contract refunds the locked asset to the sender, ensuring no funds are lost.
HTLCs provide a cryptographic mechanism for atomic payments, ensuring fairness by either transferring assets upon correct disclosure of the preimage or refunding the sender in case of timeout.
In the following section, we leverage these techniques to design our proposed scheme.
4 xRWA
This section begins with an overview of xRWA, followed by a detailed description of the proposed scheme.
4.1 Overview
Figure 2: xRWA comprises three functional layers: (i) asset identification and tokenization, where DIDs and VCs are used to represent real-world assets on-chain; (ii) cross-chain authentication, which prevents repeated registration across heterogeneous blockchains; and (iii) cross-chain interaction through channel-based settlement, enabling efficient trading of RWAs.
xRWA provides a framework for RWAs that supports secure identification, cross-chain authentication, and efficient interoperability across heterogeneous blockchains. Figure 2 shows the workflow, which consists of three layers: asset identification and tokenization, cross-chain authentication, and cross-chain interaction.
The process begins with the digital representation of off-chain assets such as real estate, stablecoins, and private credit. Traditional approaches rely on centralized custodians or special-purpose vehicles that create contractual mappings between physical assets and their digital forms. xRWA instead adopts Decentralized Identifiers (DIDs) and Verifiable Credentials (VCs) for decentralized identity management. By embedding customized attributes into VCs, the framework binds issuers and holders to asset records, ensuring legitimacy while reducing reliance on intermediaries. After tokenization, assets may need to be recognized across multiple blockchains. Existing methods often require repeated verification on each chain. xRWA introduces a cross-chain authentication protocol based on Simplified Payment Verification (SPV) proofs. By linking DIDs and VCs with tokenized assets, and applying SPV validation, authentication performed on one chain can be reused across others. Cross-chain transfers also require efficient settlement. xRWA proposes a cross-chain channel that supports settlement without channel closure. This allows frequent interactions to occur off-chain, with only final states recorded on-chain, reducing on-chain costs.
In the following, we provide a detailed description of each component of the framework, covering asset identification and tokenization, cross-chain authentication, and cross-chain interaction.
4.2 Identification and Tokenization
To realize the identification of RWAs, we construct $\mathsf{RWA\text{-}CC}$ (RWA Composite Credential) using decentralized identifiers (DIDs) and verifiable credentials (VCs). The credential follows a modular design with four components: $\mathsf{asset}$, $\mathsf{identity}$, $\mathsf{compliance}$, and $\mathsf{custody}$. Each component includes its own $\mathsf{sStatus}$ and $\mathsf{sProof}$ to support independent revocation and verification, while a top-level proof guarantees overall integrity.
The $\mathsf{asset}$ component defines the core identification and tokenization of a real-world asset. The top-level $\mathsf{id}$specifies the identifier of the credential instance itself. Within the $\mathsf{asset}$ object, the field $\mathsf{assetId}$ assigns a DID-based unique identifier to the asset. The field $\mathsf{assetType}$ indicates the broad category of the real-world asset, while $\mathsf{category}$ provides a finer classification within that type, for example distinguishing $\mathsf{Residential}$ from $\mathsf{Commercial}$ property. The $\mathsf{classDid}$ links the asset to a class definition maintained by an authoritative registry, enabling semantic grouping and interoperability across assets of the same class. A crucial element is $\mathsf{tokenBinding}$, which establishes the binding between the off-chain asset record and its tokenized representation on a blockchain. It records the token standard, the target chain, the contract address, and the token identifier in a unified structure.
The $\mathsf{identity}$ component extends the description of an asset beyond its basic identifier. The field $\mathsf{schemaVersion}$ specifies the version of the identity schema, while $\mathsf{identitySchema}$ provides a reference URI that defines its structure. The $\mathsf{identifiers}$ array records registration information from official registries; each entry contains details such as the registration scheme, the identifier value, the jurisdiction, and the issuing authority. The $\mathsf{taxonomies}$ field captures standardized classification codes together with descriptive labels, ensuring interoperability across different cataloguing systems. Spatial information is anchored by $\mathsf{spatialFootprint}$, which encodes the asset’s geographic location in GeoJSON format, including its geometry and granularity [14]. Additional evidence is attached through $\mathsf{documents}$, which may include deeds, certificates, or licenses, each linked to cryptographic hashes and media metadata. The $\mathsf{relations}$ field links the asset to other entities or classes, while $\mathsf{attributes}$ provide structured factual details such as size or physical characteristics. The $\mathsf{custom}$ object further allows local extensions, for example policy tags or IoT-based identifiers. Finally, the component includes $\mathsf{sStatus}$ for revocation or suspension tracking and $\mathsf{sProof}$ for cryptographic verification.
The $\mathsf{compliance}$ component records regulatory information that governs the circulation of the asset. The field $\mathsf{licenseId}$ identifies the license or permit under which the asset is authorized, while $\mathsf{sellableRegions}$ lists the jurisdictions where transfer is legally allowed. Restrictions such as $\mathsf{NoCrossBorderSale}$ are specified in $\mathsf{restrictions}$, and the effective period is defined by $\mathsf{effectiveFrom}$ and $\mathsf{effectiveTo}$. The $\mathsf{regulatorDid}$ provides the decentralized identifier of the authority that endorses or enforces these conditions. To ensure trustworthy usage, the component also includes $\mathsf{sStatus}$ for revocation or suspension tracking and $\mathsf{sProof}$ for independent cryptographic verification.
The $\mathsf{custody}$ component captures information related to the safekeeping of the asset. It specifies the custodian through $\mathsf{custodianDid}$, together with the physical or organizational $\mathsf{location}$ where the asset is maintained. The $\mathsf{policy}$ field encodes applicable standards such as security or compliance frameworks, and $\mathsf{auditCycleDays}$ defines the frequency of regular audits. Insurance arrangements are linked through $\mathsf{insurancePolicyRef}$. Similar to the compliance component, $\mathsf{sStatus}$ is included to manage revocation or suspension, and $\mathsf{sProof}$ ensures verifiable integrity.
After an RWA is identified through $\mathsf{RWA\text{-}CC}$, the next step is to tokenize it on a blockchain. The process builds on the algorithms for decentralized identifiers (DIDs) and verifiable credentials (VCs) introduced in Section 3. The issuer $\mathcal{I}$ generates its decentralized identity using $\mathsf{DID.KeyGen}$ and obtains a DID $did_{\mathcal{I}}$. The asset $a$ is also assigned a DID $did_{a}$, which is included in the $\mathsf{asset}$ component of $\mathsf{RWA\text{-}CC}$. The issuer then executes $\mathsf{VC.Issue}$ to generate a credential $\mathsf{cred}$ that encapsulates the attributes of $a$, including $\mathsf{identity}$, $\mathsf{compliance}$, and $\mathsf{custody}$. This binds the off-chain asset to its verifiable description under the issuer’s DID. To make the asset tradable, the credential is linked to an on-chain token via $\mathsf{tBinding}$, which specifies the token standard (e.g., ERC-721), the blockchain, the contract address, and the identifier. This ensures that smart contracts interacting with the token can invoke $\mathsf{VC.Verify}$ against the issuer’s DID and status list to check its validity. Finally, when a holder $\mathcal{U}$ presents the token in applications, only the necessary attributes are disclosed through $\mathsf{VC.Prove}$, ensuring selective transparency while maintaining compliance guarantees.
1"id": "did:ion:EiAa...credential123",
2"asset": {
3 "assetId": "did:ion:EiAa...asset123",
4 "assetType": "RealEstate",
5 "category": "Residential",
6 "classDid": "did:web:issuer.example.org:class:RE-RESIDENCE",
7 "tokenBinding": {
8 "standard": "ERC-721",
9 "chain": "eip155:1",
10 "contract": "0xBC4CA0EdA7647A8aB7C2061c2E118A18a936f13D",
11 "tokenId": "1234"
12 }
13},
14"identity": {
15 "schemaVersion": 1,
16 "identitySchema": "https://example.org/schemas/rwa-identity-v2.json",
17 "identifiers": [
18 {
19 "identifierScheme": //Identification System,
20 "identifierValue": "example-2024-000123",
21 "jurisdiction": //Jurisdiction Code,
22 "issuingAuthority": //Issuing or Registration Agency
23 }
24 ],
25 "taxonomies": [
26 {"system": "UNSPSC", "code": "70131500", "label": "Residential property"}
27 ],
28 "spatialFootprint": {
29 "encoding": "GeoJSON",
30 "geometry": {
31 "type": "Polygon",
32 "coordinates": [[[121.45,31.20],[121.46,31.20],[121.46,31.21],[121.45,31.21],[121.45,31.20]]]
33 },
34 "granularity": "parcel"
35 },
36 "documents": [
37 {"name": "Deed", "hash": "0xabc...", "mediaType": "application/pdf", "issuedBy": "did:web:issuer.org"}
38 ],
39 "relations": [
40 {"relation": "belongsToClass", "target": "did:web:issuer.example.org:class:RE-RESIDENCE"}
41 ],
42 "attributes": [
43 {"name": "floorArea", "value": "120", "unit": "sqm"}
44 ],
45 "custom": {
46 "localPolicyTag": "RWA-POL-01",
47 "iotDeviceIds": ["sensor-xyz-001"]
48 },
49 "sStatus": {...},
50 "sProof": {...}
51},
52"compliance": {
53 "licenseId": "example-2025-8899",
54 "sellableRegions": [//Jurisdiction Code],
55 "restrictions": ["NoCrossBorderSale"],
56 "effectiveFrom": "2025-01-01",
57 "effectiveTo": "2026-01-01",
58 "regulatorDid": "did:web:regulator.gov",
59 "sStatus": {...},
60 "sProof": {...}
61},
62"custody": {
63 "custodianDid": "did:web:custody.bank.example",
64 "location": "Vault-example-01",
65 "policy": "ISO-27001+SOC2",
66 "auditCycleDays": 30,
67 "insurancePolicyRef": {"hash": "0xins..."},
68 "sStatus": {...},
69 "sProof": {...}
70}
When an RWA is deployed to a single blockchain, the tokenization process described above ensures that its digital representation is securely bound to a DID and a verifiable credential. However, in practice, RWAs often need to migrate across chains. A typical scenario resembles Web2 systems where an asset record, once canceled in one registry, must be reissued in another. Similarly, an RWA token may be burned on a chain $C_{1}$ and later reissued or transferred on $C_{2}$ to access new markets. Without interoperability, the tokenization procedure has to be repeated independently on each chain, leading to redundant authentication and increased operational overhead.
To overcome this limitation, we next introduce a cross-chain authentication protocol that allows an RWA verified on one chain to be recognized on another without repeating the full tokenization workflow.
4.3 Cross-Chain Authentication
Consider an RWA asset $a$ that has been tokenized on $C_{1}$, producing a credential $\mathsf{cred}$ anchored by a transaction $tx$ in block $B$. To enable recognition of $a$ on another chain $C_{2}$ without re-executing the full tokenization procedure, we employ SPV as defined in Section 3.3.
[…]
Unlike conventional approaches where each blockchain must independently repeat the full verification of decentralized identifiers and verifiable credentials, the SPV-based protocol requires such verification only once on the source chain. Subsequent chains merely validate the compact SPV proof to ensure that the authenticated transaction has indeed been recorded on the source ledger and that the embedded commitment is consistent with the disclosed credential fields. This design eliminates redundant executions of DID resolution, credential signature checks, and revocation list queries across multiple ledgers, thereby reducing both computation and communication overhead while maintaining consistency of authentication results.
The detailed procedure of the SPV-based cross-chain authentication is summarized in Algorithm 1.
Algorithm 1 SPV-based Cross-Chain Authentication
1:Setup: Asset $a$ is tokenized on $C_{1}$, producing a credential $\mathsf{cred}$ anchored by transaction $tx$ in block $B$.
2:Commitment (on $C_{1}$): Issuer embeds into $tx$ a commitment binding the asset identifier, digest of $\mathsf{cred}$, token binding, and auxiliary parameters. txtx is included in $B$, whose header contains the Merkle root $\mu$.
3:Proof generation (off-chain): Compute $\pi\leftarrow\text{SPV}.\allowbreak\mathsf{Prove}\allowbreak(tx,B)$, yielding $\pi$ = $(p,\mu)$, where $p$ is the Merkle path authenticating txtx. Submit $(tx,\pi)$ to $C_{2}$.
4:Verification (on $C_{2}$): Verifier executes $\text{SPV}.\allowbreak\mathsf{Verify}\allowbreak(tx,\allowbreak\pi)$. If valid, it parses the commitment in txtx and checks consistency with disclosed fields of $\mathsf{cred}$, including issuer DID, revocation status, and jurisdictional restrictions.
5:Acceptance: If all checks pass, $\mathsf{cred}$ is accepted on $C_{2}$ as if locally issued.
4.4 Efficient Cross-Chain Interactions
In this part, we adopt HTLCs and off-chain channels to construct a cross-chain channel that supports partial settlement of RWAs without requiring channel closure. We first provide the definition of the cross-chain channel and then describe how it can be applied to the settlement of RWAs.
Cross-Chain Channel. A cross-chain channel [13] aims to guarantee atomicity while significantly reducing on-chain interactions. It allows parties to perform frequent cross-chain payments off-chain and only settle on-chain when necessary.
Definition 6 (Cross-Chain Channel)
The process of this channel can be represented by a tuple of polynomial-time algorithms $\Pi\overset{\text{def}}{=}(\mathsf{Open},\allowbreak\mathsf{Update},\allowbreak\mathsf{Lock},\allowbreak\mathsf{Unlock},\allowbreak\mathsf{Refund},\allowbreak\mathsf{Close})$.
  • $\mathsf{Open}$: Each party deposits assets into contracts on their respective chains, forming a bi-directional channel.
  • $\mathsf{Update}$: During the channel lifetime, the two parties exchange signed off-chain states to update balances. These updates do not require on-chain transactions, which improves efficiency.
  • $\mathsf{Lock}$: The parties submit the negotiated interaction results to the blockchain and lock their respective assets. Hash-locked conditions $h(\rho)$ are placed on both chains with timeouts $T_{1}>T_{2}$ to guarantee atomicity.
  • $\mathsf{Unlock}$: If the receiver discloses the preimage $\rho$ within the time limit $T_{2}$, it can unlock the locked assets on one chain. Once $\rho$ is revealed, the sender can use the same $\rho$ to unlock the corresponding assets on the other chain.
  • $\mathsf{Refund}$: If the preimage $\rho$ is not disclosed before the timeout, the locked asset is returned on-chain to the original depositor’s account, and the channel state is updated accordingly.
  • $\mathsf{Close}$: When the channel is terminated, the remaining deposits locked in the channel are returned on-chain to the respective parties’ accounts, thereby releasing the channel balance and concluding the channel lifecycle.
The above definition shows that settlement of RWAs can be achieved within the channel lifecycle through the $\mathsf{Unlock}$ and $\mathsf{Refund}$ procedures, without requiring channel closure. This feature allows partial settlements to be executed while keeping the channel open, thereby reducing the cost of repeated channel establishment and improving the efficiency of cross-chain interactions.
Cross-Chain Interactions. First, the parties invoke $\Pi.\mathsf{Open}$ by deploying contracts $\xi^{\mathsf{1}}$ and $\xi^{\mathsf{2}}$ on $C_{1}$ and $C_{2}$, which establish the channel environment and record deposits. During the lifetime of the channel, they repeatedly apply $\Pi.\mathsf{Update}$ to exchange signed off-chain states describing tentative transfers of different subsets of RWAs. For example, a state $\sigma_{j}$ may specify that $\mathcal{B}$ pays $v_{j}$ units on $C_{1}$ in exchange for batch $S_{j}\subseteq\{a_{1},\dots,a_{m}\}$. These updates remain entirely off-chain, enabling multiple negotiations without incurring blockchain overhead. Once the parties agree on a settlement, $\Pi.\mathsf{Lock}$ is executed: both sides upload the agreed final state to $\xi^{\mathsf{1}}$ and $\xi^{\mathsf{2}}$, where hash-locked conditions $h(\rho)$ with timeouts $T_{1}>T_{2}$ are installed to ensure atomicity. If $\mathcal{B}$ reveals the preimage $\rho$ on $\xi^{\mathsf{2}}$ before $T_{2}$, $\Pi.\mathsf{Unlock}$ delivers the agreed union of RWAs $\bigcup_{j}S_{j}$ to $\mathcal{B}$, and allows $\mathcal{S}$ to redeem the corresponding aggregated payment on $\xi^{\mathsf{1}}$ before $T_{1}$. If instead $\mathcal{B}$ does not reveal $\rho$ in time, $\Pi.\mathsf{Refund}$ is triggered: $\xi^{\mathsf{2}}$ returns the RWAs to $\mathcal{S}$ at $T_{2}$, and $\xi^{\mathsf{1}}$ returns the funds to $\mathcal{B}$ after $T_{1}$. Finally, once the session concludes, $\Pi.\mathsf{Close}$ is called to release any remaining deposits and terminate the channel.
Through this process, multiple RWAs can be traded within the channel via successive $\mathsf{Update}$ operations, while only a single round of on-chain $\mathsf{Unlock}$ or $\mathsf{Refund}$ is needed to settle the results. This reduces on-chain congestion and lowers the overall settlement costs. The detailed procedure of the cross-chain channel for multiple RWAs is summarized in Algorithm 2.
Algorithm 2 Efficient Cross-Chain Interactions
1:Participants: Buyer $\mathcal{B}$ (funds on $C_{1}$), Seller $\mathcal{S}$ (RWAs $\{a_{1},\dots,a_{m}\}$ on $C_{2}$).
2:Contracts: $\xi^{\mathsf{1}}$ on $C_{1}$, $\xi^{\mathsf{2}}$ on $C_{2}$.
3:Parameters: deposits $d_{\mathcal{B}},d_{\mathcal{S}}$, preimage $\rho$, hash $h(\cdot)$, timeouts $T_{1}>T_{2}$.
4:Channel Opening (on-chain):
[…]
6:Off-chain Negotiation (loop):
[…]
8:On-chain Locking (commit final state):
[…]
10:Settlement (on-chain):
11:if $\mathcal{B}$ reveals $\rho$ before $T_{2}$ then
12:  $\Pi.\mathsf{Unlock}$ on $\xi^{\mathsf{2}}$: deliver $\bigcup_{j}S_{j}$ to $\mathcal{B}$.
13:  $\Pi.\mathsf{Unlock}$ on $\xi^{\mathsf{1}}$: release the aggregated payment to $\mathcal{S}$ before $T_{1}$.
14:else
15:  $\Pi.\mathsf{Refund}$ on $\xi^{\mathsf{2}}$: return $\{a_{1},\dots,a_{m}\}$ to $\mathcal{S}$ at $T_{2}$.
16:  $\Pi.\mathsf{Refund}$ on $\xi^{\mathsf{1}}$: return $d_{\mathcal{B}}$ (and any unsettled funds) to $\mathcal{B}$ at $T_{1}$.
17:end if
18:Channel Continuation / Closure: The channel may remain open for further $\Pi.\mathsf{Update}$ rounds without reinstalling locks. Optionally, $\Pi.\mathsf{Close}$ releases remaining deposits and terminates the channel.
In the following section, we demonstrate the feasibility of our scheme by conducting simulation-based experiments.
5 Implementation and Performance Evaluation
In this section, we present our concrete implementation of xRWA and test its performance.
5.1 Implementation
Experiments are conducted on a workstation with an Intel Core i9-13980HX CPU and 32 GB RAM, running Windows 11 (24H2). Python 3.12.10 is used to implement and benchmark the VC routines and the SPV algorithm, and the reported execution times are obtained from these Python implementations under multi-threaded settings. Smart-contract experiments are performed on Remix VM (Prague) using Solidity 0.8.30 to measure gas consumption for the HTLC and cross-chain channel workflows. The entire codebase consists of more than 2,000 lines, including the Python implementations for VC/SPV and the Solidity contracts for on-chain evaluation. Our code repository is publicly available at https://github.com/ZhuHaoming2005/xRWA.
5.2 Performance Evaluation
Table 1 presents the sizes (in KB) of VCs across different asset types. The covered categories include vehicles, real estate (RE), gold, art, bonds, funds, and intellectual property (IP). The sizes range from 6.95 KB to 8.05 KB, with an overall average of 7.27 KB.
Table 1: The size of VC across different RWA types.
Type Size (KB)
Vehicle 7.11
RE 8.05
Gold 6.95
Art 7.38
Bond 7.09
Fund 7.27
IP 7.10
Average 7.27
Figure 3 reports the performance of issuance and verification under a multi-threaded workload (500 credentials, 100 iterations, 8 worker threads). The average issuance latency is 8.16ms per VC, with a P95 latency of 9.04ms. Verification is faster, with an average latency of 0.96ms per VC and a P95 of 1.27ms. These results show that credential issuance requires on the order of several milliseconds per VC, while verification can be performed in less than one millisecond on average. The measured latencies indicate that both operations are lightweight and can be executed efficiently under multi-threaded workloads.
Figure 3: Latency in issuing and verifying VCs.
Figure 4 reports the empirical validation of the $\mathcal{O}(\log n)$ complexity of SPV verification. The measured latency follows the fitted logarithmic model, where $n$ denotes the number of transactions per block.
$t_{\text{SPV}}(n)\approx 0.69\times\log_{2}(n)+0.23\ \text{$\mu$s}$
[…]
Figure 4: Verification time of SPV proofs under varying transaction counts.
For the cross-chain channel, handling $n$ cross-chain interactions requires only 917,253 gas in total, since all updates are performed off-chain and only the opening and unlocking operations are executed on-chain. In contrast, the HTLC scheme consumes 465,426 $\times\,n$ gas for $n$ interactions, as each interaction must execute both the lock and unlock operations on-chain. This demonstrates that the cross-chain channel achieves significantly higher efficiency than HTLC, especially when the number of interactions increases.
6 Conclusion
This paper presents a cross-chain framework for RWAs that addresses identification, authentication, and interaction in a unified workflow. The design leverages DIDs and VCs for decentralization, applies an SPV-based protocol to prevent redundant verification, and introduces a cross-chain channel that enables partial settlement without closure. Experiments indicate that the framework reduces on-chain costs and improves efficiency in cross-chain settings. As future work, we explore potential security vulnerabilities in RWA projects and develop mitigation strategies to strengthen their reliability in practice.
[…]